SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (31st July, 2007)

Attendees: Debbie Bard, Kim Hojeong, Su Dong, Mark Tibbetts.

Operations

  • Pages last night for repeated synch requests from DCZ. ODF ROM log files showed repeated "MPIC: spurious interrupt" messages. I told the pilot to power cycle the crates and that fixed it.
  • Saturday AM: ODF damage ended a run, with most systems giving white alarms. DCZ was the only one which did not clear and did not allow the pilot to start a new run (luckily, PEP was doing a little work and there were no beams expected). The ODF ROM log files showed the message: " ERROR : conflict for NEW TSF-Y map says 5555 while I detect 404" which I hadn't seen before, and which meant the TSF boards had not all been detected. Re-booting the ROMs did nothing, so I had the pilot power-cycle the crate which solved the problem.
    Later, it was heard from ODF that the timing signal and fiducials had been lost from PEP, which ended the run. The TSF crates always need to be rebooted when this happens.
  • Hojoeng will start training next week for being trigger ops manager.
  • Debbie (and Hojeong) will be out of town during the last week of August, so someone will need to cover the pager then...

EMT

  • Mark has been looking at TrigTools package, and has had some problems having it work out of the box, related to reading the FCT word.

Trigger Configuration

  • Hojeong has been looking into re-doing the trigger statistics study with SP9 MC. Except for one case (where the MC was used with no further cuts applied) the analysts who first performed the study need to help out. Only Selina has responded that she can re-do the study, David needs to request the MC still and no response from the others.
  • Will there be a new L3 filter produced for the 1E line? We might have to tighten the other lines if we keep on with the high luminosity and add the 1E, to make the rate manageable.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.